The standards for one case of liquid weed killer are as follows: Direct materials 5 lb @ $ 6.80 /lb Direct labor 2.4 hr @ $ 14.00 /hr Variable overhead (based on machine hours) 1.5 hr @ $ 4.50 /hr During the week ended May 6, the following activity took place: 2,910 machine hours were worked. 11,400 lb of raw material were purchased for inventory at a total cost of $80,940. 1,900 cases of finished product were produced. 9,260 lb of raw material were used. 4,420 labor hours were worked at an average rate of $14.35 per hour. $12,513 actual variable overhead costs were incurred. Required: Calculate each of the following variances Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

a. Price variance for raw materials purchased.   
b. Raw materials usage variance.
c. Direct labor rate variance.
d. Direct labor efficiency variance.
e. Variable overhead spending variance.
f. Variable overhead efficiency variance.
Answer #1

a) Price variance = (Standard price-actual price)*Actual qty = (6.8*11400-80940) = 3420 U

b) Usage variance = (standard qty-actual qty)Standard price = (1900*5-9260)*6.8 = 1632 F

c) Labor rate variance = (Standard rate-actual rate)actual hours = (14-14.35)*4420 = 1547 U

d) Labor efficiency variance = (Standard hour-actual hour)Standard rate = (1900*2.4-4420)*14 = 1960 F

e) Variable overhead spending variance = (Standard rate-actual rate)actual hours = (4.5*2910-12513) = 582 F

f) Variable overhead efficiency variance = (Standard hour-actual hour)Standard rate = (1900*1.5-2910)*4.5 = 270 U